OpenStudy (anonymous):

Solve y over negative 4 + 5 = 15

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{y}{-4}+5=15 }\) subtract 5 from both sides, \(\Huge\color{blue}{ \sf \frac{y}{-4}+5 \color{red} { -5 } =15 \color{red} { -5 } }\) and you get, \(\Huge\color{blue}{ \sf \frac{y}{-4}=15 }\) now, let's multiply both sides times -4, \(\Huge\color{blue}{ \sf \frac{y}{-4} \color{red} { \times (-4) } =15 \color{red} { \times (-4) } }\) and you get \(\Huge\color{blue}{ \sf y=-60 }\)

OpenStudy (anonymous):

Alright, thanks. If i have anymore questions ill @ you into it. Big help bro.

OpenStudy (solomonzelman):

I'll try whenever I can ;)

OpenStudy (anonymous):

Okay but, the only answers are A.80 B.-80 C.40 D.-40 @SolomonZelman

OpenStudy (solomonzelman):

I forgot to subtract 5 from 15, it would be y/(-4) = 10 y= -40

OpenStudy (anonymous):

Okay, Thanks bro.

OpenStudy (solomonzelman):

You welcome

OpenStudy (anonymous):

Solve for x: 3 − (2x − 5) < −4(x + 2) @SolomonZelman @AnswerMyQuestions